| Name | Email | Date | Comment
|
| ANS
| anolla@naic.edu
| 6/7/2000
| Receiver Tsys. it's too high. Christ Salter try to use the system last night
but it had a high tsys. (it was not in scheduled but systems has to be ready at any time)
|
| ANS
| anolla@naic.edu
| 8/17/2000
| The antanna was removed from the receiver in order to check it at 700
Mhz(L.W. is working on it)
|
| ANS
| anolla@naic.edu
| 8/18/2000
| The 611 feed was tested for possible operation at 700-724 Mhz .Seems
good ! 611 feed replaced at 3:30 P.M. by L.W. and E. Gallo .
Did some total power observing with this recvr. and also some correlator
output. It is working OK.
|
| ANS
| anolla@naic.edu
| 8/29/2000
| Replaced 610-612 filter in chn B post amp chassis with a 700-703 mhz
filter for evaluation.
|
| ANS
| anolla@naic.edu
| 8/30/2000
| Birdie at 702.24 mhz , audio FM from TV channel 54 , this is not where it
is originally broadcast at, which is 715.75 (difference of 13.5 mhz)
Tried to do cals today on chn B at 700 mhz,but the 702 birdies is back. 610-612 MHz
filter replaced in chn B.
|
| ANS
| anolla@naic.edu
| 12/28/2000
| JR installed the receiver in the new open frame cabinet.
|
| ANS
| anolla@naic.edu
| 5/3/2001
| L Wray removed the 611 mhz filters in order to modify the chassis to add the
611/700 mhz filter bank.
|
| ANS
| anolla@naic.edu
| 5/4/2001
| Receiver installed back after modification to add the 611/700 mhz filter bank.
Channel A Tsys was 122 ºk, channel B was 107 ºk.
|
| ANS
| anolla@naic.edu
| 5/11/2001
| L Wray & EG installed the 611/&00 Mhz filter bank on 5/10/2001.They tested it
& looks to be working fine.Filter bank is at 611 mhz.
|
| ANS
| anolla@naic.edu
| 5/23/2001
| The A/L sw at the radar room bypassed , recvr always on antanna except when
430 MHz xmmter on CH 1 up.
|
| ANS
| anolla@naic.edu
| 7/2/2001
| System was recalibrated on 6/28/2001 for the 700 Mhz operation , its values are :
Chann. A 86.1 ºk Chann. B 62.6 ºk
|
| ANS
| anolla@naic.edu
| 7/9/2001
| Receiver taken down to replace the directional couplers.The new couplers are
MAC model C320230 , sn 320001 is on chnn A and sn 320002 is on B . Miteq RF
amps are model AMF-3F-005010-06-L , sn 416968 chnn A and sn 427580 on B.
|
| ANS
| anolla@naic.edu
| 8/ /2001
| System installed upstairs with new D.C. for lower cal ( 30 db couplers).
|
| ANS
| anolla@naic.edu
| 8/22/2001
| System was wrongly wired for A/L protection on the SB xmitter mode by ANS
(me). Was rewired wrigth , now.
|
| ANS
| anolla@naic.edu
| 8/30/2001
| L Wray calibrated the recvr , it looks to be +/- 36 ºk each chann. She used both
diodes to recheck , the 15 db en and the 5 db en.
|
| ANS
| anolla@naic.edu
| 9/20/2001
| Chann A first stage amplifier was replaced with the spare one , it had a
Tsys of 300ºk, instead its normal one (around 150ºk).
|
| ANS
| anolla@naic.edu
| 10/8/2001
| The recvr got burned out again (same chnn A ) , Genesh brought it down on
10/4/2001.
|
| ANS
| anolla@naic.edu
| 10/24/2001
| 611 mhz A/L sw installed thru the CB/SBW shutter control relay in order to
protect it from the SB xmitter.
|
| ANS
| anolla@naic.edu
| 5/31/2002
| 611 mhz system was installed yesterday direct into the 611 mhz filters ( temp
controlled) bypassing the 611/700 mhz filter bank. Also there are new R.F. amps on it
these are Angle Linear amps.Its description is as follow:
Output compression level are typically +18 dBm and 3rd order intercept is typically +29
dBm. Higher output level units are
available, see our "HY" series of PHEMT preamplifiers.
Noise figure is typically 0.5 dB below 500 MHz, 0.6 dB to 750 MHz and 0.7 db to 1
GHz.
Gain is typically: 37 db at 300 MHz; 34 db at 400 MHz; 32 dB at 500 MHz; 30 db at 600
MHz; 27 dB at 800 MHz; and 26 dB at
950 MHz. Input and output return loss are typically greater than 10 dB.
|
| ANS
| anolla@naic.edu
| 6/5/2002
| LW calibrated the 611 Mhz receiver using the noise figure meter.
|
| ANS
| anolla@naic.edu
| 8/5/2002
| 611 Mhz filter installed back , was been retuned to 608 mhz at the edge of its low
side of the passband.Temperature for chan 1 is 87.4 ºf and 86.6 ºf for chan B.
|
| LSW | lwray@naic.edu | Wednesday, August 7, 2002 15:21:47 | Re-installed filters, after being mechanically re-tuned to lower frequencies (608 - 612). Tsys same at 130K
|
| ans | anolla@naic.edu | Monday, August 26, 2002 10:04:47 | Receiver chann A was not working properly ,LWR found a faulty cable on it.Back to service.
|
| ans | anolla@naic.edu | Friday, August 30, 2002 11:46:23 | The old antanna was installed back (bazooka match) ,hybrid one down at the shop.
|
| ans | anolla@naic.edu | Monday, October 20, 2003 14:04:05 | The new shutter driver at the Dome was producing acomb generated rfi to all low frequency receivers at the Dome(327,430,611,etc.)
|
| ans | anolla@naic.edu | Tuesday, December 30, 2003 13:31:15 | System taken down by Ed Castro.
|
| ans | anolla@naic.edu | Wednesday, January 7, 2004 16:15:29 | System removed to a new location at the rotary floor.
|
| lsw | lwray@naic.edu | Friday, January 16, 2004 09:53:40 | receiver hooked back up on Jan 15th.
|
| lsw | lwray@naic.edu | Monday, February 2, 2004 15:53:32 | Cal not firing since Thursday due to control card coming out of socket in the 611 Power Supply chassis. This card was reattached and a screw put in to hold in place.
Couldn't check with Tsys, althought Ant/Load and cals are ok.
|
| ans | anolla@naic.edu | Wednesday, June 30, 2004 11:20:35 | All rf connectors were check and cleaned.Also rf feedthrus tighted at the chassis.
|
| lsw | lwray@naic.edu | Thursday, July 1, 2004 15:08:00 | Calixto and Carlos did tests on this receiver, looking at RFI after the filter.
|
| ans | anolla@naic.edu | Monday, August 30, 2004 13:10:08 | System removed from upstairs in order to be checked. Tsys is a bit higher on chan. B.
|
| ans | anolla@naic.edu | Friday, September 3, 2004 15:53:34 | System back to normal after a check out. Tsys on chann A is 130 K , B is 121 K.
|
| ans | anolla@naic.edu | Wednesday, January 12, 2005 13:11:37 | Add an AC line cord guard to the recvr. chassis.Also to the recvr shutter control P.S. & to Ivan Cordero shutter box.
|
| lsw | lwray@naic.edu | Thursday, January 27, 2005 12:48:09 | Jan 25,26: Testing front end, took spectrum analyzer plots to show the RFI present in the 700-740 MHz band.
|
| ans | anolla@naic.edu | Tuesday, May 17, 2005 12:58:33 | Rey Velez & Calixto were searching for rfi into this band.RFI is causing problems to the 611 Mhz channel A receiver.It was postponed due to factor time.430 Mhz xmitter test on schedule.
|
| cheo | jrosa@naic.edu | Wednesday, June 22, 2005 08:37:52 | on 21 june 05 found rf output cable A lose.
|
| ans | anolla@naic.edu | Friday, March 30, 2007 11:34:20 | The 611 Mhz system retired from service.
|